OpenStudy (anonymous):

In a bag containing only black and green balloons, 6 balloons are black. If the probability of choosing a black balloon from the bag is 1/3, how many green balloons are in the bag?

OpenStudy (anonymous):

let X be total balloons prob of getting black balloon is 6/X=1/3 X=18, the total ballons now green balloons= 18-6=12

OpenStudy (anonymous):

Probability of getting black is 1/3, so probability of getting green is 2/3. If 1/3=6 2/3=?